CEA
HPC Calcul haute performance
AccueilComplexe de calculLogiciels pour le HPCRecherche et developpementsCollaborations

 

 

IBM

 

Oracle

 

DDN

 

Bull
Home / HPSS User Forum 2012


HPSS User Forum 2012

Agenda

CEA
HPSS


Monday October 15th, 2012

 

 

8h30

Registration and breakfast

9h30

Welcome and presentation of agenda

9h50

HPSS Executive Commitee Update

10h20

Break

10h35

Delivery process, plans, policies (Thuy Tran)

11h35

HPSS release and development plans (Paul Montag)

12h35

Lunch

14h10 Summary Release and Feature Review (Dave Fisher)

14h40

Burning issues process, review of 2011 burning issues

15h15 Site status: Indiana University (Kurt Seiffert)
15h30 Site status: NCAR (Bill Anderson)

15h45

Break

16h

Vendor presentation : DDN

17h Evolving HPSS, including Panda (David Boomer)
17h30 Site status: ORNL (Jason Anderson)
17h45 Site status: MPI fuer Plasmaphysics (Manuel Panea-Doblado)

18h

End of session

 


Hotel Pullman Bercy

 

 

Tuesday October 16th, 2012

 

 

8h30

Breakfast

9h

HPSS Deployment and Testing Details (Dennis Tran)

9h30 DB2 for HPSS update (David Boomer)
10h05 Site status : KEK (Lida Yoshimi)

10h20

Break

10h35

GridFTP with HPSS at NCSA (Jason Alt)

11h05 Oracle Storage: a complete portfolio to store,
protect and preserve mass data storage

12h10

Lunch

13h45

Transfer by bus to CEA/TGCC

15h

HPC data storage at CEA: past & future (J.C Lafoucrière)

15h30 Lustre-HSM project update (J.C Lafoucrière)
15h55 Break

16h10

Site status: ECMWF (Francis Dequenne)

16h25

Migration to Linux (Francis Dequenne)

16h55 Presentation of TGCC and Curie supercomputer

17h05

Visit of CEA/TGCC computing center

18h30

Transfer by bus to the Abbaye des Vaux-de-Cernay

19h30

Cocktail & dinner at the Abbaye des Vaux-de-Cernay

23h

Transfer by bus to Hotel Pullman, Paris

 


Vaux de Cernay

 

 

Wednesday October 17th, 2012

 

 

8h30

Breakfast

9h

Use of HPSS's VFS Layer (Craig Tierney)

9h30

Site status: DWD

9h45

Site status: Met Office (Mark Francis)

10h

Met Office - Recent Challenges of Expanding Our Current Tape Libraries

10h30

Break

10h45

UDAs (David Boomer)

11h15 HPSS Training plans (Kayla Broussard)
11h45 Site status: Scinet (Jaime Pinto)

12h00

Lunch

13h45

HPSS RAIT at NCSA (Jason Alt)

14h15 Gatekeeper/Accounting
14h45 Site status: LANL (M'hamed Jebbanema)
15h Site status: IN2P3 (Pierre-Emmanuel Brinette)
15h15 Open discussion

15h45

Break

16h

Repack (Dave Fisher)

16h30 Site status: LLNL (Todd Heer)
16h45 Site status: NOAA/RDHPCS (Craig Tierney)
17h Burning issues 2012

18h

Break before social event

19h30

Appointment for social event

 

Social event: dinner-cruise on Seine River

 


Melody Blues Peniche

 

 

Thursday October 18th, 2012

 

 

8h30

Breakfast

9h

IBM GPFS 3.5 (Antoine Tabary)

10h GHI update (Kayla Broussard)

10h30

Break

10h45

Open discussion

12h

Wrap-up, closing remarks, next HUF, …

12h30

* End of HUF2012 *



HUF 2012 is sponsored by :
IBM      Oracle       DDN       Bull